OpenStudy (anonymous):

Find the angle between the given vectors, round to the nearest tenth degree u= i -j , v=3i + 5j

OpenStudy (amistre64):

use the property that:\[|u||v|\cos\alpha=u\cdot v\]

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

its 14 degrees right?